# Semiclassical Gravity Beyond General Relativity: Insights from Torsion

### v1

@@ 5.3.2 Renormalization ambiguities
$\alpha_{8}$ $\begin{aligned} &-2\Big\{-g_{(a|m}(U_{|b)\ell cdef}+U_{\ell|b)cdef})\\ &\times K^{(\ell|cd}K^{|m)ef}\\ &+U^{(f\ell)}_{\quad\ (a|cde}(K_{f|b)}^{\quad c}-K_{f\ |b)}^{\ \ c})K_{\ell}^{\ de}\\ &+U^{(f\ell)}_{\quad\ cd(a|e}K_{f}^{\ cd}(K_{\ell|b)}^{\quad e}-K_{\ell\ |b)}^{\ e})\\ &-\frac{1}{2}U^{(\ell m)}_{\quad\ \ cdef}K_{\ell}^{\ cd}K_{m}^{\ \ ef}g_{ab}\Big\}\end{aligned}$ $6(U_{(cf)}^{\quad\ abde}-U^{b\ \ a\ \ de}_{\ (c\ \ f)})K^{f}_{\ de}$
$\Theta_{ab}$ $\Sigma_{c}^{\ ab}$
$\beta_{1}$ $\begin{aligned} &-4\{(2K^{f}_{\ (a|d}-K^{f}_{d(a|})K_{ecf}\epsilon_{|b)}^{\quad cde}\\ &-K^{\ell}_{\ ce}K_{fd\ell}\epsilon^{cdef}g_{ab}\}\end{aligned}$ $2\{\epsilon^{adb}_{\quad e}K^{e}_{\ dc}+\epsilon^{ad\ e}_{\ \ c}K^{b}_{\ de}\}$
$\beta_{2}$ $\begin{aligned} &-2\{\mathring{G}_{ab}+K^{d}_{\ (ab)}K^{c}_{\ cd}-K^{d}_{\ c(a}K^{c}_{\ b)d}\\ &-K^{d\ \ c}_{\ [c}K^{e}_{\ e]d}g_{ab}\}\end{aligned}$ $\{g^{ab}K^{d}_{\ dc}-\delta^{a}_{\ c}K^{d\ b}_{\ d}+K_{c}^{\ ba}-K^{b\ a}_{\ c}\}$
@@
From inspecting Eq. (5.3.2), it follows that $\beta_{2}$ renormalizes Newton’s constant, while $\gamma$ plays the role of a cosmological constant (which is known to have relevant applications in GR [41, 42]). Furthermore, in semiclassical GR with a nonminimally coupled Klein–Gordon field, only the terms associated with $\alpha_{1},\ \alpha_{2},\ \beta_{2}$, and $\gamma$ contribute [32]. This is consistent with our findings, since the only terms with a nontrivial dependence on $K^{c}_{\ ab}$ are those coupled to these constants, along with $\beta_{1}$ and $\alpha_{3}$. However, the $\beta_{1}$ term is the Holst term, which vanishes in GR, whereas the term coupled to $\alpha_{3}$ is a boundary term in GR. In particular, the Hodge dual of Eq. (100) is proportional to $R^{\mu\nu}_{\ \ \alpha\beta}R_{\mu\nu\gamma\delta}\epsilon^{\alpha\beta\gamma\delta}$, while the dual of the term coupled to $\tilde{\alpha}_{3}$ is $R_{\alpha\beta}^{\ \ \ \mu\nu}R_{\gamma\delta\mu\nu}\epsilon^{\alpha\beta\gamma\delta}$. In GR, these expressions coincide; however, this equivalence no longer holds in the presence of torsion.
Lastly, substituting Eqs. (82) and (86) into the equations of motion reveals that the semiclassical equations differ significantly from their classical counterparts. In the metric equation, both scaling and renormalization ambiguities give rise to fourth-order derivative terms, which complicate the analysis of the Cauchy problem [43, 44, 45, 46, 47]. In addition, semiclassically, the contorsion equation ceases to be purely algebraic. Consequently, in a semiclassical treatment, it may occur that $K^{c}_{\ ab}\neq 0$ in regions where $\omega(\sigma_{c}^{\ ab})=0$. A detailed analysis of this result is left for future work.
@@ 5.4 Conformal anomaly
Lastly, we compute $\omega(\tau^{a}_{\ a})$, ignoring renormalization ambiguities, for simplicity, to compare it with the classical expression $\tau^{a}_{\ a}$. From Eqs. (60), (62), (86), and (89), and for the parameters that make the classical theory conformally invariant, $m=0$ and $\xi=1/6$, we find
$\omega(\tau^{a}_{\ a})=\frac{1}{4\pi^{2}}\,\mathring{\nabla}_{c}\!\left\{\frac{1}{3}g^{ab}K^{c}_{\ ab}[w-v_{0}\ln(M^{2})]\right\}+\frac{v_{1}}{4\pi^{2}}.$ (106)
@@ 6 Conclusions
In this work, we show that the Hadamard renormalization procedure for a nonminimally coupled, free, massive Klein–Gordon field can be extended to spacetimes with torsion, at least within Einstein–Cartan theory. The axiomatic framework we propose is a direct generalization of Wald’s axioms and proves sufficient to establish a well-defined renormalization scheme. Using Hadamard states, we regularize $\omega(\tau_{ab})$ and $\omega(\sigma_{c}^{\ ab})$ by subtracting their divergence structure. This subtraction naturally introduces scale and renormalization ambiguities, which we can explicitly compute.
@@
Moreover, the semiclassical theory provides a qualitatively different physical description compared to the classical one. In Einstein–Cartan theory, the equation of motion for $K^{c}_{\ ab}$ is algebraic, so this tensor is only nonvanishing in regions where $\sigma_{c}^{\ ab}\neq 0$. However, the renormalization ambiguities modify the contorsion equation, turning it into a differential rather than an algebraic equation. It remains an open question whether this feature allows for the detection of torsion outside polarized matter. Still, one of the main takeaways of this work is that semiclassical effects provide access to a rich phenomenology within modified gravity theories that could otherwise remain inaccessible. Therefore, the semiclassical framework serves as a valuable avenue for exploring such theories.
@@ Acknowledgments
7 Appendix: Geometrical bitensors
@@
$g_{b}^{\ b^{\prime}}\,\mathring{\nabla}_{b^{\prime}}\mathring{\nabla}_{a}\sigma$ $=-g_{ab}-\frac{1}{6}\,\mathring{R}_{ac_{1}bc_{2}}\,\mathring{\nabla}^{c_{1}}\sigma\,\mathring{\nabla}^{c_{2}}\sigma+\mathcal{O}(\sigma^{3/2}),$ (109)
where unprimed indices are associated with $x$ and primed indices with $x^{\prime}$. In addition, $g_{a}^{\ a^{\prime}}$ is the bitensor that parallel-transports vectors from $x$ to $x^{\prime}$ along the geodesic connecting them. This bitensor satisfies
@@
Further details on $g_{a}^{\ a^{\prime}}$ can be found in Refs. [30, 32].
@@
$+\frac{1}{72}\mathring{R}\mathring{R}_{ab}-\frac{1}{30}\mathring{R}_{a}^{\ c}\mathring{R}_{bc}+\frac{1}{60}\mathring{R}^{cd}\mathring{R}_{acbd}$
$+\frac{1}{60}\mathring{R}_{a}^{\ cde}\mathring{R}_{bcde}\Big)\mathring{\nabla}^{a}\sigma\mathring{\nabla}^{b}\sigma+\mathcal{O}(\sigma^{3/2}),$ (119)

### v2

@@ 5.3.2 Renormalization ambiguities
$\alpha_{8}$ $\begin{aligned} &-2\Big\{-g_{(a|m}(U_{|b)\ell cdef}+U_{\ell|b)cdef})K^{(\ell|cd}K^{|m)ef}\\ &\qquad+U^{(f\ell)}_{\phantom{(f\ell)}\ (a|cde}(K_{f|b)}^{\phantom{f|b)}c}-K_{f\phantom{c}|b)}^{\phantom{f}c})K_{\ell}^{\phantom{\ell}de}\\ &\qquad+U^{(f\ell)}_{\phantom{(f\ell)}\ cd(a|e}K_{f}^{\phantom{f}cd}(K_{\ell|b)}^{\phantom{\ell|b)}e}-K_{\ell\phantom{e}|b)}^{\phantom{\ell}e})\\ &\qquad-\frac{1}{2}U^{(\ell m)}_{\phantom{(\ell m)}cdef}K_{\ell}^{\phantom{\ell}cd}K_{m}^{\phantom{m}ef}g_{ab}\Big\}\end{aligned}$ $6(U_{(cf)}^{\phantom{(cf)}abde}-U^{b\phantom{(c}a\phantom{f)}de}_{\phantom{b}(c\phantom{a}f)})K^{f}_{\phantom{f}de}$
$\Theta_{ab}$ $\sigma_{c}^{\phantom{c}ab}$
$\beta_{1}$ $4\{(2K^{f}_{\phantom{f}(a|d}-K^{f}_{\phantom{F}d(a|})K_{ecf}\epsilon_{|b)}^{\phantom{|b)}cde}-K^{\ell}_{\phantom{\ell}ce}K_{fd\ell}\epsilon^{cdef}g_{ab}\}$ $2\{\epsilon^{adb}_{\phantom{adb}e}K^{e}_{\phantom{e}dc}+\epsilon^{ad\phantom{c}e}_{\phantom{ad}c}K^{b}_{\phantom{b}de}\}$
$\beta_{2}$ $2\{\mathring{G}_{ab}+K^{d}_{\phantom{d}(ab)}K^{c}_{\phantom{c}cd}-K^{d}_{\phantom{d}c(a}K^{c}_{\phantom{c}b)d}-K^{d\phantom{[c}c}_{\phantom{d}[c}K^{e}_{\phantom{e}e]d}g_{ab}\}$ $\{g^{ab}K^{d}_{\phantom{d}dc}-\delta^{a}_{\phantom{a}c}K^{d\phantom{d}b}_{\phantom{d}d}+K_{c}^{\phantom{c}ba}-K^{b\phantom{c}a}_{\phantom{b}c}\}$
@@
From inspecting Eq. (5.3.2), it follows that $\beta_{2}$ renormalizes Newton’s constant, while $\gamma$ plays the role of a cosmological constant (which is known to have relevant applications in GR [41, 42]). Furthermore, in semiclassical GR with a nonminimally coupled Klein–Gordon field, only the terms associated with $\alpha_{1},\ \alpha_{2},\ \beta_{2}$, and $\gamma$ contribute [32]. This is consistent with our findings, since the only terms with a nontrivial dependence on $K^{c}_{\phantom{c}ab}$ are those coupled to these constants, along with $\beta_{1}$ and $\alpha_{3}$. However, the $\beta_{1}$ term is the Holst term, which vanishes in GR, whereas the term coupled to $\alpha_{3}$ is a boundary term in GR. In particular, the Hodge dual of Eq. (100) is proportional to $R^{\mu\nu}_{\phantom{\mu\nu}\alpha\beta}R_{\mu\nu\gamma\delta}\epsilon^{\alpha\beta\gamma\delta}$, while the dual of the term coupled to $\tilde{\alpha}_{3}$ is $R_{\alpha\beta}^{\phantom{\alpha\beta}\mu\nu}R_{\gamma\delta\mu\nu}\epsilon^{\alpha\beta\gamma\delta}$. In GR, these expressions coincide; however, this equivalence no longer holds in the presence of torsion.
Lastly, substituting Eqs. (82) and (86) into the equations of motion reveals that the semiclassical equations differ significantly from their classical counterparts. In the metric equation, both scaling and renormalization ambiguities give rise to fourth-order derivative terms, which complicate the analysis of the Cauchy problem [43, 44, 45, 46, 47]. In addition, semiclassically, the contorsion equation ceases to be purely algebraic. Consequently, in a semiclassical treatment, it may occur that $K^{c}_{\phantom{c}ab}\neq 0$ in regions where $\omega(\sigma_{c}^{\phantom{c}ab})=0$. A detailed analysis of this result is left for future work.
@@ 5.4 Conformal anomaly
Lastly, we compute $\omega(\tau^{a}_{\phantom{a}a})$, ignoring renormalization ambiguities, for simplicity, to compare it with the classical expression $\tau^{a}_{\phantom{a}a}$. From Eqs. (60), (62), (86), and (89), and for the parameters that make the classical theory conformally invariant, $m=0$ and $\xi=1/6$, we find
$\omega(\tau^{a}_{\phantom{a}a})=\frac{1}{4\pi^{2}}\,\mathring{\nabla}_{c}\!\left\{\frac{1}{3}g^{ab}K^{c}_{\phantom{c}ab}[w-v_{0}\ln(M^{2})]\right\}+\frac{v_{1}}{4\pi^{2}}.$ (106)
@@ 6 Conclusions
In this work, we show that the Hadamard renormalization procedure for a nonminimally coupled, free, massive Klein–Gordon field can be extended to spacetimes with torsion, at least within Einstein–Cartan theory. The axiomatic framework we propose is a direct generalization of Wald’s axioms and proves sufficient to establish a well-defined renormalization scheme. Using Hadamard states, we regularize $\omega(\tau_{ab})$ and $\omega(\sigma_{c}^{\phantom{c}ab})$ by subtracting their divergence structure. This subtraction naturally introduces scale and renormalization ambiguities which do not depend on the state and can be explicitly computed.
@@
Moreover, the semiclassical theory provides a qualitatively different physical description compared to the classical one. In Einstein–Cartan theory, the equation of motion for $K^{c}_{\phantom{c}ab}$ is algebraic, so this tensor is only nonvanishing in regions where $\sigma_{c}^{\phantom{c}ab}\neq 0$. However, the renormalization ambiguities modify the contorsion equation, turning it into a differential rather than an algebraic equation. It remains an open question whether this feature allows for the detection of torsion outside polarized matter. Still, one of the main takeaways of this work is that semiclassical effects provide access to a rich phenomenology within modified gravity theories that could otherwise remain inaccessible. Therefore, the semiclassical framework serves as a valuable avenue for exploring such theories.
@@ Acknowledgments
Appendix A Geometrical bitensors
@@
$g_{b}^{\phantom{b}b^{\prime}}\,\mathring{\nabla}_{b^{\prime}}\mathring{\nabla}_{a}\sigma$ $=-g_{ab}-\frac{1}{6}\,\mathring{R}_{ac_{1}bc_{2}}\,\mathring{\nabla}^{c_{1}}\sigma\,\mathring{\nabla}^{c_{2}}\sigma+\mathcal{O}(\sigma^{3/2}),$ (109)
where unprimed indices are associated with $x$ and primed indices with $x^{\prime}$. In addition, $g_{a}^{\phantom{a}a^{\prime}}$ is the bitensor that parallel-transports vectors from $x$ to $x^{\prime}$ along the geodesic connecting them. This bitensor satisfies
@@
Further details on $g_{a}^{\phantom{a}a^{\prime}}$ can be found in Refs. [30, 32].
@@
$+\frac{1}{72}\mathring{R}\mathring{R}_{ab}-\frac{1}{30}\mathring{R}_{a}^{\phantom{a}c}\mathring{R}_{bc}+\frac{1}{60}\mathring{R}^{cd}\mathring{R}_{acbd}$
$+\frac{1}{60}\mathring{R}_{a}^{\phantom{a}cde}\mathring{R}_{bcde}\Big)\mathring{\nabla}^{a}\sigma\mathring{\nabla}^{b}\sigma+\mathcal{O}(\sigma^{3/2}),$ (119)